A fixed mass of gas of volume 600 cm3 at a temperature of 27° c is cooled at constant pressure to a temperature of 0° c. What is the change in volume
1
Expert's answer
2020-04-15T10:13:15-0400

"\frac{V}{T}=const"

"\frac{V_1}{T_1}=\frac{V_2}{T_2}",

were "V_1" and "V_2" - initial and final volume of the gas, respectively; "T_1" and "T_2" - initial and final temperature of the gas, respectively.

"V_2=\frac{V_1 \cdot T_2}{T_1}"

"T [K] \approx T[ \degree C] +273"

"T_2 = 27 +273 = 300 \; K"

"T_1 = 0 +273 = 273 \; K"

"V_2=\frac{600 \cdot 273}{300} = 546 \; cm^3"

"\Delta V=V_2-V_1"

"\Delta V=546-600=-54 \; cm^3"


Answer: volume of the gas decrease by 54 cm3.